Dr. D.B. Carpenter Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S018 1GS

April 12, 1994

Council Tax Billing Section, Civic Centre, Southampton, S09 4XD

Dear Council Tax Billing Section,

I have been forwarded a council tax bill for the 94-95 period, directed to my former address—Top Flat, 98 Gordon Avenue, Southampton S02 1DA.

I moved to the address quoted above in February. Please could you update your records and advise me of any possible changes to my rate of taxation.

Yours sincerely,